High Resolution Laser Spectroscopy of the C̃^{2}A_{1}–X̃^{2}A_{1} Transition of SrNH_{2} and B̃^{2}B_{1}–X̃^{2}A_{1} Transition of CaNH_{2}

\underline{Chunfeng Zhao}, J.W. Hepburn, P.F.Bernath, Center for Molecular Beams and Laser Chemistry, University of Waterloo, Waterloo, Ont.Canada, N2L 3G1.

The high-resolution laser excitation spectrum of C̃^{2}A_{1}–X̃^{2}A_{1} transition of SrNH_{2} and B̃^{2}B_{1}–X̃^{2}A_{1} transition of CaNH_{2} were recorded. The SrNH_{2} and CaNH_{2} molecules were made by pulsed laser ablation of Sr and Ca metal rods followed by reaction with NH_{3}. A cw dye laser was used to detect the laser induced fluorescence from a supersonic beam. The measured line positions have been fitted to give molecular constants for the electronic states.